OpenStudy (ggdvaaa):

A printer printed 68 letters in 4 seconds. If it prints letters at a constant speed, it can print ______ letters in 30 seconds.

jabez177 (jabez177):

68 letters in 4 seconds so first we need to find how many letters in 1 second. Still remember how to do that? :)

OpenStudy (ggdvaaa):

think i forgot

OpenStudy (ggdvaaa):

;D

jabez177 (jabez177):

letters / time

jabez177 (jabez177):

68 divided by 4

jabez177 (jabez177):

68 divided by 4 is what? :)

OpenStudy (anonymous):

You can use a proportion so \[4/68 = 30/x\]

OpenStudy (anonymous):

68*30 = 2040

OpenStudy (anonymous):

2040/4 = 510

OpenStudy (ggdvaaa):

17 sorry i was on another tab

OpenStudy (anonymous):

510 would be the answer

jabez177 (jabez177):

Ohhh yeah... I forgot it's the other way around... @unthinkablejoe is right.

OpenStudy (anonymous):

:p

OpenStudy (ggdvaaa):

ohhh

jabez177 (jabez177):

But yes, 17 is the answer. Now they want to know how many letters it can print in 30 seconds. So we multiply 17 by 30 and that will get you your answer. :)

OpenStudy (ggdvaaa):

wait so lol 17 is the answer? i thought it was 510

OpenStudy (anonymous):

510 is the answer

OpenStudy (anonymous):

good job

jabez177 (jabez177):

510

OpenStudy (ggdvaaa):

ohh ok thank youu both :))

jabez177 (jabez177):

When I said 17 is the answer, I meant the thing can type 17 letters in 1 second. :P

OpenStudy (ggdvaaa):

ohh ok xD